# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ V-B Feasibility of the Disturbed Bang-Singular-Bang Trajectory
Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$. Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$, and then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{\left(p\right)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{\left(p\right)}}\right)$ is continuous in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$. If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i^{\left(p\right)}_{j-1}}\right)+d_{p}<0$, then let $r_{j}^{\left(p\right)}=0$; otherwise, according to Proposition 3, Proposition 4, and Proposition 5, $\exists 1\leq r_{j}^{\left(p\right)}\leq n$, s.t. $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$. Let $t_{1,j}^{\left(p\right)}>t_{i^{\left(p\right)}_{j-1}}$ is small enough, s.t. $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)\leq\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+1$ for $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$. $t_{2,j}^{\left(p\right)}<t_{i^{\left(p\right)}_{j}}$ is constructed similarly. Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}\right)$. In both case, $\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ due to continuity. Let
$\varepsilon<-\frac{\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}}{2C\left\|{\boldsymbol{c}}_{p}\right\|}.$ (39)
Then, $\forall t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$,
${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}<0.$ (40)
In other words, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible in $\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$.
The feasibility of ${\boldsymbol{x}}^{\prime}$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)\cup\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ needs further discussion. Let
$\begin{dcases}\varepsilon<t_{1,j}^{\left(p\right)}-t_{i^{\left(p\right)}_{j-1}},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}+1\right)\right)^{-1},\\ \varepsilon<\left(C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}+1}+1\right)\right)^{-1},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+8\right)^{-1}.\end{dcases}$ (41)
Denote $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}$. Then, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right]$,
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)$
$=$ $\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
$\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}-\varepsilon,t_{i^{\left(p\right)}_{j-1}}\right]$, it has
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (43)
$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)+1}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$
$\leq$ $\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2.$
@@
Therefore, if $t_{i^{\left(p\right)}_{j-1}}^{\prime}\geq t_{i^{\left(p\right)}_{j-1}}$, then $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$. If $t_{i^{\left(p\right)}_{j-1}}^{\prime}<t_{i^{\left(p\right)}_{j-1}}$, then (30b) implies that $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\leq\frac{3}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$. $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j-1}}\right]$,
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (44)
$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)+\left(\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2\right)\varepsilon$
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$.
@@
Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)$. Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$. Similarly, let $\varepsilon$ is small enough, and it has ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ in $\left[t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right]$.
@@
In summary, when $\varepsilon$ is small enough, $\forall p\in\left[P\right]$, $t\in\left[t_{0}^{\prime},t_{M}^{\prime}\right]$, it has ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ holds.
@@
Thirdly, the inequality $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is small enough. ∎
@@
Theorem 2 points out that if the disturbed trajectory is represented by the same augmented switching law to the original trajectory and the disturbance is light enough, then the disturbed trajectory is feasible. In other words, Theorem 2 provides a novel variational approach under fixed augmented switching laws with a feasibility guarantee near the constrained boundary.
V-C Necessary Condition for the Optimal Trajectory

### v2

@@ V-B Feasibility of the Perturbed BBS Trajectory
Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$ and $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$. Then, $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ is continuous in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$. If $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ holds, then let $r_{j}^{(p)}=0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$; otherwise, according to Propositions 4, 5, and 6, $\exists r_{j}^{(p)}\in\left[n\right]$, s.t. $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ and $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$. Let $t_{1,j}^{(p)}\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$ is sufficiently small, s.t. $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ holds $\forall t\in\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$. $t_{2,j}^{(p)}\leq t_{i^{(p)}_{j}}$ is constructed similarly. Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{(p)},t_{i^{(p)}_{j}}\right)$.
Due to the continuity of $f_{p}$, $\sup_{t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]}f_{p}\left(t\right)<0$ holds. For the part far from keypoints, i.e., $\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$, let
$\varepsilon<-\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\right)^{-1}\sup\nolimits_{t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]}f_{p}\left(t\right).$ (38)
Then, according to (30), $\forall t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$, it holds that
$f_{p}^{\prime}\left(t\right)\leq f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}<0.$ (39)
In other words, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$.
Consider the part near the left keypoint, i.e., $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$. Assume that $t_{i^{(p)}_{j-1}}^{\prime}<t_{1,j}^{(p)}$; otherwise, $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]=\varnothing$.
For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$, let $\varepsilon$ be sufficiently small, s.t. $f_{p}\left(t\right)\leq\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$. By (30) and (38), $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$, $f_{p}^{\prime}\left(t\right)\leq 0$ is feasible since
$f_{p}^{\prime}\left(t\right)\leq$ $f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (40)
$\leq$ $\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)-\frac{1}{4}f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0.$
Figure 3: Feasibility of the perturbed trajectory near the keypoints. (a) Constrained arc. (b) Unconstrained arc.
For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)=0$ and $t_{i_{j-1}^{(p)}}<t_{1,j}^{(p)}$, let
$\varepsilon<\hat{\varepsilon}\triangleq-\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}+1\right)^{-1}.$ (41)
Hence, $\forall t\in\left[t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right]$, it holds that
$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
If $t_{i^{(p)}_{j-1}}^{\prime}<t_{i^{(p)}_{j-1}}$, let $\varepsilon$ be sufficiently small, s.t.
$\sup\nolimits_{\delta\in\left[0,\varepsilon\right]}\left\|{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}-\delta\right)-{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)\right\|<\hat{\varepsilon}.$ (43)
@@
Then, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]\subset\left[t_{i^{(p)}_{j-1}}-\varepsilon,t_{i^{(p)}_{j-1}}\right]$, it holds that
$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r_{j}^{(p)}-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ (44)
$+$ $\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left(\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)\right\|\right)$
$\leq$ $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
@@
Therefore, once $\varepsilon$ is sufficiently small, then $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}<0$ holds in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$. ${\mathcal{L}}^{\prime}={\mathcal{L}}$ implies that $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right)$. Therefore, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible. Similarly, once $\varepsilon$ is sufficiently small, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible in $\left[t_{2,j}^{(p)},t_{i^{(p)}_{j}}^{\prime}\right]$.
@@
Note that the above conditions on $\varepsilon$ only depends on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$. In summary, once $\varepsilon$ is sufficiently small, then $\forall p\in\left[P\right]$, $t\in\left[t_{0}^{\prime},t_{M}^{\prime}\right]$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ holds.
@@
Thirdly, the control constraints $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is sufficiently small. ∎
@@
Theorem 2 points out that if the perturbed trajectory is represented by the same ASL to the original trajectory and the perturbance is small enough, then the perturbed trajectory is feasible. In other words, Theorem 2 provides a novel variational approach under the fixed ASL with a feasibility guarantee near constraints’ boundaries.
V-C State-Centric Necessary Condition of Optimality